OpenStudy (toxicsugar22):

#1 the final burret reading mL is 31.12 initial buret reading mL IS 15.72 volume of titrant mL is 15.49 Moles of NaOH used is 1.568*10^-3 mol moles of HC2H3O2 neutralized by the NaOH is 1.568*10^-3 mole molarity of dil HC2H3O2 is .0784 M #2 final buret reading is 15.53 initial buret reading mL is 0.00 volume of titrant mL is 15.53 Moles of NaOH used is 1.57*10^-3 Moles of HC2H302 neutralized by the NaOH is 1.57*10^-3 molarity of dil HC2H3O2 is .0785 M

OpenStudy (toxicsugar22):

how did we find the moles of HC2H302

OpenStudy (triciaal):

what exactly is the question?

OpenStudy (triciaal):

are you suppose to check which of the 4 choices is true?

OpenStudy (triciaal):

I am not sure what you are asking

OpenStudy (triciaal):



OpenStudy (triciaal):

moles of HC2H3O2= 1.568*10^-3 mole M1 molecular mass of HC2H3O2 = 1 + (1282) +3+(16*2) = molarity of dil HC2H3O2 is .0784 M = moles/volume incomplete

OpenStudy (triciaal):

incomplete means not finished not complete there should be more

OpenStudy (triciaal):

the volume is 15.49 i think this would be V1

OpenStudy (triciaal):

m2 = Molarity * molecular mass* volume

OpenStudy (triciaal):

what is the molecular mass for error above molecular mass of HC2H3O2 = 1 + (12*2) +3+(16*2) = 60

OpenStudy (triciaal):

let me try and get you some help

OpenStudy (triciaal):

well read some of the post above

OpenStudy (triciaal):

@cutepochacco can you please help to clarify

OpenStudy (triciaal):

@jim_thompson5910, @satellite73 can you or do you know who can help in Chem?

OpenStudy (triciaal):

Hero said "Definitely not me. I'm not a fan of chemicals. However, she can try her luck on Slader" Jim said "sorry I don't know chem or anyone who is good at it maybe the chem section has someone?"

OpenStudy (triciaal):

did you find m1 and v1 above?

OpenStudy (triciaal):

m1 = 1.568 * 10^-3 * 60.06 v1 = 15.72ml v2 = 31.12 ?

OpenStudy (anonymous):

triciaal is right nd defined whole question clearly so plz read carefully

OpenStudy (triciaal):

moles = molarity * volume

OpenStudy (triciaal):

ml to liter * 10^-3 yes and this was already stated

OpenStudy (triciaal):

@aaronq will you please review and advise sorry this is so long please help

OpenStudy (triciaal):

as stated before what exactly is the question? what you are calculating is already given

OpenStudy (triciaal):

volume used is 20ml = 0.02L molarity = moles = volume * molarity

OpenStudy (triciaal):

@Phebe can you help please?

OpenStudy (triciaal):

I tried earlier to get you help. I don't remember much of this stuff not fresh on my mind. I want you to have the right steps so you can learn what to do. Find someone in Chemistry. you can tag me for Math.

OpenStudy (toxicsugar22):

molarity of standardized NaOH solution is 0.1012 M Volume of diluted vinegar used in each titration 20 mL 0.2 l

OpenStudy (anonymous):

ehh, sorry not fresh enough on my mind. would have to do some reading.

OpenStudy (anonymous):

M*V=M'*V'

OpenStudy (toxicsugar22):

yes ,m1v1=m2v2

OpenStudy (toxicsugar22):

so what is m1 v1 m2 v2 in the problem

OpenStudy (toxicsugar22):

that i am confused about

OpenStudy (triciaal):

do not get stressed or discouraged. chemistry is really not hard. lot to understand yes but it's not recent enough for me to help you properly. I would love to help but I would have to review the material and I really don't have time to do that. you also need to read slowly to absorb more.

OpenStudy (toxicsugar22):

is v1 = 15.72ml is v2 = 31.12

OpenStudy (anonymous):

Create a table: x (x-mean) (x-mean)^2 .0784 .0785 .0789 .07895 Add up the last column and plug it into the standard deviation formula

OpenStudy (toxicsugar22):

.0784 .0785 .0789 .07895 ADDED TOGETHER is 0786875 M

OpenStudy (toxicsugar22):

I need to find the Moles of HC2H302 neutralized by the NaOH?

OpenStudy (toxicsugar22):

the final burret reading mL is 31.12 initial buret reading mL IS 15.72 volume of titrant mL is 15.49 Moles of NaOH used is 1.568*10^-3 mol

OpenStudy (toxicsugar22):

i FOUND THIS Moles of NaOH used is 1.568*10^-3 mol

OpenStudy (toxicsugar22):

Now I need to find what the Moles of HC2H302 neutralized by the NaOH?

OpenStudy (arabpride):

Havn't even taken chemistry - so sorry, -wish i could help ,-,

OpenStudy (anonymous):

what do you want to calculate?

OpenStudy (toxicsugar22):

@kellyconnelly

OpenStudy (harsha19111999):

Hey. You can calculate it by using the formula M1V1 = M2V2

OpenStudy (toxicsugar22):

how